OpenStudy (anonymous):

Evaluate the following sums sum (2k)/(k-1), i=2 to 5

OpenStudy (anonymous):

Can you list all steps for understanding?

OpenStudy (anonymous):

for i=2, (2*2)/(2-1) = ??? for i=3, (2*3)/(3-1) = ??? for i=4, (2*4)/(4-1) = ??? for i=5, (2*5)/(5-1) = ??? now just add them up....

OpenStudy (anonymous):

wait... is the "k" supposed to be "i" ???

OpenStudy (anonymous):

the I is supposed to be a K sorry typo

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw... :)